# Anosov representations, strongly convex cocompact groups and weak eigenvalue gaps

### v3

@@ 3. The contraction property
[…]
$\mathsf{T}_{hP_{1}^{+}}\mathbb{P}(\mathbb{R}^{d})$ =$=\Big{\{}dL_{h}d\pi^{+}\left(X\right):X\in\bigoplus_{i=2}^{d}\mathbb{R}E_{i1}\Big{\}}$
$\mathsf{T}_{hP_{1}^{-}}\mathsf{Gr}_{d-1}(\mathbb{R}^{d})$ =$=\Big{\{}dL_{h}d\pi^{-}\left(X\right):X\in\bigoplus_{i=2}^{d}\mathbb{R}E_{1i}\Big{\}}.$
For $u\in\{0\}\times\mathbb{R}^{d-1}$ we denote by $X_{u}^{+}\in\mathsf{T}_{hP_{1}^{+}}\mathbb{P}(\mathbb{R}^{d})$ and $X_{u}^{-}\in\mathsf{T}_{hP_{1}^{-}}\mathsf{Gr}_{d-1}(\mathbb{R}^{d})$ the tangent vectors
@@
in the fibers of the bundles $\sigma_{\ast}\mathcal{E}^{\pm}\rightarrow\Gamma\backslash\hat{\Gamma}$ over $x=[\hat{m}]_{\Gamma}$ and $\pi^{+},\pi^{-}$ are the projections from $\mathsf{SL}_{d}(\mathbb{R})$ to $\mathbb{P}(\mathbb{R}^{d})$ and $\mathsf{Gr}_{d-1}(\mathbb{R}^{d})$ respectively.
The following lemma shows that when the geodesic flow on $\sigma_{\ast}\mathcal{E}^{-}$ is weakly contracting, then the geodesic flow on $\sigma_{\ast}\mathcal{E}^{+}$ is weakly dilating.
@@
For two sequences of positive real numbers $(a_{n})_{n\in\mathbb{N}}$, $(b_{n})_{n\in\mathbb{N}}$ we write $a_{n}\asymp b_{n}$ if there exists $R>0$ such that $R^{-1}a_{n}\leqslant b_{n}\leqslant Ra_{n}$ for every $n\in\mathbb{N}$. We may assume that $\rho(\Gamma)$ is contained in $\mathsf{SL}^{\pm}_{d}(\mathbb{R})$, otherwise we replace $\rho$ with $\hat{\rho}(\gamma)=|\textup{det}(\rho(\gamma))|^{-1/d}\rho(\gamma)$, $\gamma\in\Gamma$.
Let $(t_{n})_{n\in\mathbb{R}}$ be an increasing unbounded sequence. For each $n\in\mathbb{N}$, we may choose $\gamma_{n}\in\Gamma$ such that $\gamma_{n}\varphi_{t_{n}}(\hat{m})$ lies in the compact fundamental domain $\mathcal{F}$. There exist $k_{1n},k_{2n}\in K$ so that
@@
Notice that for $g\in P_{1}^{\pm}$ we have $dL_{g}\circ d\pi^{\pm}=d\pi^{\pm}\circ\textup{Ad}(g)$. An elementary calculation shows that
@@
$=dL_{k_{1n}}\left(d\pi^{+}\left(\begin{pmatrix}0&0\\ \frac{1}{\lambda_{n}}A_{n}u&0\end{pmatrix}\right)\right)$
and similarly we check that
@@
$\left\|\varphi_{t_{n}}\big{(}X_{u}^{+}\big{)}\right\|_{\varphi_{t_{n}}(x)}\asymp\frac{\left\|A_{n}u\right\|}{|\lambda_{n}|}\ \ \mathrm{and}\ \ \left\|\varphi_{t_{n}}\big{(}X_{u}^{-}\big{)}\right\|_{\varphi_{t_{n}}(x)}\asymp|s_{n}|\left\|B_{n}^{-t}u\right\|,$
where |||||$\cdot$||| | ⋅ | | denotes the usual Euclidean norm on $\mathbb{R}^{d-1}$. Up to passing to a subsequence, we may assume that $\lim_{n}\gamma_{n}\varphi_{t_{n}}(\hat{m})=\hat{m}^{\prime}$. Since the maps $\tau^{\pm}$ are continuous we conclude, up to subsequence, that $(\gamma_{n}\tau^{+}(\hat{m}))_{n\in\mathbb{N}}$ and $(\gamma_{n}\tau^{-}(\hat{m}))_{n\in\mathbb{N}}$ converge to $\tau^{+}(\hat{m}^{\prime})\in\partial_{\infty}\Gamma$ and $\tau^{-}(\hat{m}^{\prime})\in\partial_{\infty}\Gamma$ respectively. We have $\xi^{+}(\tau^{+}(\gamma_{n}\hat{m}))=k_{1n}P_{1}^{+}$ and $\xi^{-}(\tau^{+}(\gamma_{n}\hat{m}))=k_{2n}P_{1}^{-}$ and by transversality, there exist $p_{n}\in P_{1}^{+},q_{n}\in P_{1}^{-}$ and $g\in G$ such that $\lim_{n}k_{1n}p_{n}=\lim_{n}k_{2n}q_{n}=g$. Then there exist $z_{n},z_{n}^{\prime}\in\mathbb{R}$ so that $\lim_{n}z_{n}k_{1n}e_{1}=ge_{1}$ and $\lim_{n}z_{n}^{\prime}k_{2n}e_{1}=g^{-t}e_{1}$ and we observe that $|z_{n}|,|z_{n}^{\prime}|$ converge respectively to $||ge_{1}||$ and $||g^{-t}e_{1}||$. Notice that $\lim_{n}z_{n}z_{n}^{\prime}\langle k_{1n}e_{1},k_{2n}e_{1}\rangle=|\langle g^{-t}e_{1},ge_{1}\rangle|=1$ and so $\lim_{n}\langle k_{1n}e_{1},k_{2n}e_{1}\rangle=\frac{1}{||ge_{1}||\cdot||g^{-t}e_{1}||}$. Recall that $k_{2n}^{-1}k_{1n}\begin{pmatrix}[0.8]\lambda_{n}&\ast\\ 0&A_{n}\end{pmatrix}=\begin{pmatrix}[0.8]s_{n}&0\\ \ast&B_{n}\end{pmatrix}$, hence by looking at the $(1,1)$ entry of both sides, we obtain ${\left|\frac{s_{n}}{\lambda_{n}}\right|=\left|\left\langle k_{1n}e_{1},k_{2n}e_{1}\right\rangle\right|}$ and so $L:=\inf_{n\in\mathbb{N}}\left|\frac{s_{n}}{\lambda_{n}}\right|>0$. Furthermore, we observe that $\begin{pmatrix}[0.6]\lambda_{n}&0\\ \ast&A_{n}^{t}\end{pmatrix}k_{1n}^{t}=\begin{pmatrix}[0.7]s_{n}&\ast\\ 0&B_{n}^{t}\end{pmatrix}k_{2n}^{t}$ and hence ${\begin{pmatrix}[0.7]\ast&\ast\\ \ast&B_{n}^{-t}A_{n}^{t}\end{pmatrix}=k_{2n}^{-1}k_{1n}}$ since $k_{1n}k_{1n}^{t}=k_{2n}k_{2n}^{t}=I_{d}$. Up to passing to a subsequence, we may assume that ${\lim_{n}B_{n}^{-t}A_{n}^{t}}=Q$ exists. Since $|\lambda_{n}||\det(A_{n})|=|s_{n}||\det(B_{n})|$ we have ${\left|\det\left(B_{n}^{-t}A_{n}^{t}\right)\right|=\left|\frac{s_{n}}{\lambda_{n}}\right|}\geqslant L>0$. In particular, $Q$ is invertible and there is $M>0$ with $\frac{1}{M}\leqslant\max(||B_{n}^{-t}A_{n}^{t}||,||A_{n}^{-t}B_{n}^{t}||)\leqslant M$ for every $n\in\mathbb{N}$. Therefore, for every $n\in\mathbb{N}$ we have
${\frac{\left\|A_{n}u\right\|}{|\lambda_{n}|}\geqslant\frac{\left\|u\right\|^{2}}{|\lambda_{n}|\left\|A_{n}^{-t}u\right\|}=\frac{\left\|u\right\|^{2}}{|\lambda_{n}|\left\|A_{n}^{-t}B_{n}^{t}(B_{n}^{-t}u)\right\|}\geqslant\frac{\left\|u\right\|^{2}}{|\lambda_{n}|\left|\left|A_{n}^{-t}B_{n}^{t}\right|\right|\cdot\left\|B_{n}^{-t}u\right\|}\geqslant\frac{L\left\|u\right\|^{2}}{M|s_{n}|\left\|B_{n}^{-t}u\right\|}},$
since $||A_{n}u||\cdot||A_{n}^{-t}u||\geqslant||u||^{2}$. Finally,
@@
(i) $\underset{n\rightarrow\infty}{\lim}||\varphi_{t_{n}}\big{(}X_{u}^{+}\big{)}||_{\varphi_{t_{n}}(x)}=+\infty$ if and only if
@@
(ii) $\underset{n\rightarrow\infty}{\lim}||\varphi_{t_{n}}\big{(}X_{u}^{-}\big{)}||_{\varphi_{t_{n}}(x)}=0$ if and only if
@@
where $\xi^{+}(x)=hP_{1}^{+}$ and $hu\in\xi^{-}(y)$. Up to passing to subsequence, we may assume that $\lim_{n}\gamma_{r_{n}}\varphi_{t_{r_{n}}}(\hat{m})$ exists and so $\lim_{n}\gamma_{r_{n}}\tau^{+}(\hat{m})\neq\lim_{n}\gamma_{r_{n}}\tau^{-}(\hat{m})$. The maps $\xi^{+}$ and $\xi^{-}$ are transverse, hence there exists $g\in G$ and $p_{n}\in P_{1}^{+}$, $q_{n}\in P_{1}^{-}$ such that $\lim_{n}\rho(\gamma_{r_{n}})hp_{n}=\lim_{n}\rho(\gamma_{r_{n}})hq_{n}=g$. Let $v_{\infty}\in e_{1}^{\perp}$ be a limit point of the sequence $\Big{(}\frac{q_{n}^{-1}u}{||q_{n}^{-1}u||}\Big{)}_{n\in\mathbb{N}}$. Then, $\lim_{n}\frac{1}{||q_{n}^{-1}u||}\rho(\gamma_{r_{n}})hu=gv_{\infty}$ and hence $\lim_{n}\sin\measuredangle\big{(}\rho(\gamma_{r_{n}})he_{1},\rho(\gamma_{r_{n}})hu\big{)}=\sin\measuredangle\big{(}gv_{\infty},ge_{1}\big{)}>0$. Since we started with an arbitrary subsequence, there exists $\varepsilon>0$ with $\big{|}\sin\measuredangle\big{(}\rho(\gamma_{r_{n}})he_{1},\rho(\gamma_{r_{n}})hu\big{)}\big{|}\geqslant\varepsilon$ for every $n\in\mathbb{N}$. Therefore, $\frac{\left\|A_{n}u\right\|}{|\lambda_{n}|}\asymp\frac{\left\|\rho(\gamma_{n})hu\right\|}{\left\|\rho(\gamma_{n})he_{1}\right\|}$. By Proposition 3.1 we have that
$\big{|}\big{|}\varphi_{t_{n}}(X_{u}^{+})\big{|}\big{|}_{\varphi_{t_{n}}(x)}\asymp\frac{\left\|A_{n}u\right\|}{|\lambda_{n}|},\ \ n\rightarrow\infty,$
and so part (i) follows. The argument for part (ii) is similar. ∎
@@ 4. The Cartan property and the uniform gap summation property
Let $G$ be a real semisimple Lie group of non-compact type, $K$ a maximal compact subgroup of $G$ and $\mu:G\rightarrow\overline{\mathfrak{a}}^{+}$ be the Cartan projection. The restricted Weyl group of $\mathfrak{a}$ in $\mathfrak{g}$ is the group $W=N_{K}(\mathfrak{a})/Z_{K}(\mathfrak{a})$, where $N_{K}(\mathfrak{a})$ (resp. $Z_{K}(\mathfrak{a})$) is the normalizer (resp. centralizer) of $\mathfrak{a}$ in $K$. The group $W$ is finite, acts simply transitively on the set of Weyl chambers of $\mathfrak{a}$ and contains a unique order two element $wZ_{K}(\mathfrak{a})\in W$ such that $\textup{Ad}(w)\overline{\mathfrak{a}}^{+}=-\overline{\mathfrak{a}}^{+}$. The element $w\in K$ defines an involution ${}^{\star}:\Delta\rightarrow\Delta$ on the set of simple restricted roots $\Delta$ as follows: if $\alpha\in\Delta$ then $\alpha^{\star}\in\Delta$ is the unique root with $\alpha^{\star}(H)=-\alpha(\textup{Ad}(w)H)$ for every $H\in\mathfrak{a}$.
@@
$\lim_{|\gamma|_{\Gamma}\rightarrow\infty}\big{\langle}\alpha,\mu(\rho(\gamma))\big{\rangle}=+\infty$
@@
$\Xi_{\theta}^{+}(g):=k_{g}P_{\theta}^{+}\ \ \textup{and}\ \ \Xi_{\theta}^{-}(g):=k_{g}wP_{\theta}^{-}.$
@@
(1) Suppose that $\rho$ admits a continuous $\rho$-equivariant map $\xi^{+}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$. The map $\xi^{+}$ satisfies the Cartan property if for any $x\in\partial_{\infty}\Gamma$ and every infinite sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of elements of $\Gamma$ with $\lim_{n}\gamma_{n}=x$,
@@
(2) Suppose that $\rho$ admits a continuous $\rho$-equivariant map $\xi^{-}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{-}$. The map $\xi^{-}$ satisfies the Cartan property if the map $\xi^{\ast}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta^{\star}}^{+}$ satisfies the Cartan property. In other words, for every $x\in\partial_{\infty}\Gamma$ and every infinite sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of elements of $\Gamma$ with $\lim_{n}\gamma_{n}=x$,
$\xi^{-}(x)=\lim_{n\rightarrow\infty}\Xi^{-}_{\theta}\big{(}\rho(\gamma_{n})\big{)}$
Remark 4.2.
Let $\rho:\Gamma\rightarrow G$ be a $P_{\theta}$-divergent representation. The Cartan property for a continuous $\rho$-equivariant map $\xi^{+}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ (resp. $\xi^{-}$) is independent of the choice of the Cartan decomposition of $G$. This follows by the fact that all Cartan subspaces of $G$ are conjugate under the adjoint action of $G$ and the second part of [22, Cor. 5.9].
@@
Suppose that $\rho,\Gamma,G$ and $\theta$ are defined as in Definition 4.1 and let $\xi:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ be a continuous $\rho$-equivariant map. Suppose that $\tau:G\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ is an irreducible $\theta$-compatible proximal representation as in Proposition 2.4 so that $\tau(P_{\theta}^{+})$ stabilizes a line in $\mathbb{R}^{d}$ and induces a $\tau$-equivariant embedding $\iota^{+}:G/P_{\theta}^{+}\xhookrightarrow{}\mathbb{P}(\mathbb{R}^{d})$. The map $\xi$ satisfies the Cartan property if and only if $\iota^{+}\circ\xi$ satisfies the Cartan property.
We need the following estimates which help us verify, in several cases, the Cartan property of limit maps into the homogeneous spaces $G/P_{\theta}^{+}$ and $G/P_{\theta}^{-}$. The second part of the following proposition has been established in [7, Lem. A4] and [22, Lem. 5.8 (i)] but for completeness we give a short proof.
@@
Let $G$ be a real semisimple Lie group, $\theta\subset\Delta$ a subset of simple restricted roots of $G$ and $\tau:G\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ an irreducible, $\theta$-proximal representation such that $\tau(P_{\theta}^{+})$ stabilizes the line $[e_{1}]$ in $\mathbb{R}^{d}$. Let $\chi_{\tau}\in\mathfrak{a}^{\ast}$ be the highest weight of $\tau$ and $g,r\in G$.
(i) If $g$ is $P_{\theta}$-proximal in $G/P_{\theta}^{+}$ with attracting fixed point $x_{g}^{+}\in G/P_{\theta}^{+}$, then
$d_{G/P_{\theta}^{+}}\big{(}x_{g}^{+},\Xi_{\theta}^{+}(g)\big{)}$ $\leqslant\exp\Big{(}-\min_{\alpha\in\theta}\big{\langle}\alpha,\mu(g)\big{\rangle}+\big{\langle}\chi_{\tau},\mu(g)-\lambda(g)\big{\rangle}\Big{)}$
(ii) If $\min_{\alpha\in\theta}\langle\alpha,\mu(g)\rangle>0$ and $\min_{\alpha\in\theta}\langle\alpha,\mu(gr)\rangle>0$, then
$d_{G/P_{\theta}^{+}}\big{(}\Xi_{\theta}^{+}(gr),\Xi_{\theta}^{+}(g)\big{)}\leqslant C_{d,r}\exp\Big{(}-\min_{\alpha\in\theta}\big{\langle}\alpha,\mu(g)\big{\rangle}\Big{)}$
where ${C_{d,r}=\sigma_{1}(\tau(r))\sigma_{1}(\tau(r^{-1}))\sqrt{d-1}}$.
@@
$g=h\begin{pmatrix}\ell^{\prime}_{1}(g)&\\ &A_{g}\\ \end{pmatrix}h^{-1}=k_{g}\exp(\mu(g))k_{g}^{\prime},\ \left|\ell_{1}^{\prime}(g)\right|=\ell_{1}(g).$
@@
$h\begin{pmatrix}\ell_{1}^{\prime}(g)&\\ &A_{g}\end{pmatrix}h^{-1}=w_{1}\begin{pmatrix}\ell_{1}^{\prime}(g)&\ast\\ &\ast\end{pmatrix}w_{1}^{-1}$
hence ${k_{g}^{-1}w_{1}\begin{pmatrix}[0.7]\ell_{1}^{\prime}(g)&\ast\\ &\ast\end{pmatrix}=}$ ${\exp(\mu(g))k_{g}^{\prime}w_{1}}$ and $\left|\left\langle k_{g}^{-1}w_{1}e_{1},e_{i}\right\rangle\right|=\frac{\sigma_{i}(g)}{\ell_{1}(g)}\left|\left\langle k_{g}^{\prime}w_{1}e_{1},e_{i}\right\rangle\right|$ for $i>1$. Therefore,
@@
Since $\min_{\alpha\in\theta}\langle\alpha,\mu(\rho(\gamma))\rangle=\log\frac{\sigma_{1}(\tau(g))}{\sigma_{2}(\tau(g))}$, $\langle\chi_{\tau},\mu(g)\rangle=\log\sigma_{1}(\tau(g))$ and $\langle\chi_{\tau},\lambda(g)\rangle=\log\ell_{1}(\tau(g))$ for $g\in G$, part (i) follows.

### v4

@@ 3. The contraction property
Let $\Gamma$ be a word hyperbolic group. Fix $\big(\hat{\Gamma},\varphi_{t}\big)$ a flow space on which $\Gamma$ acts properly discontinuously and cocompactly. Fix also $\mathcal{F}\subset\hat{\Gamma}$ a compact subset of $\hat{\Gamma}$ whose $\Gamma$-translates cover $\hat{\Gamma}$. Let $\rho:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ be a representation admitting a pair of transverse, $\rho$-equivariant maps $\xi^{+}:\partial_{\infty}\Gamma\rightarrow\mathbb{P}(\mathbb{R}^{d})$ and $\xi^{-}:\partial_{\infty}\Gamma\rightarrow\mathsf{Gr}_{d-1}(\mathbb{R}^{d})$ defining the flat section $\sigma:\Gamma\backslash\hat{\Gamma}\rightarrow\mathcal{X}_{\rho}$ of the fiber bundle $\pi:\mathcal{X}_{\rho}\rightarrow\Gamma\backslash\hat{\Gamma}$. We fix an equivariant family of norms $\big(||\cdot||_{x}\big)_{x\in\Gamma\backslash\hat{\Gamma}}$ on the fibers of the bundle $\pi_{\pm}:\mathcal{E}_{\rho}^{\pm}\rightarrow\Gamma\backslash\hat{\Gamma}$. Recall also the maps $\tau^{\pm}:\hat{\Gamma}\rightarrow\partial_{\infty}\Gamma$ defined in Subsection 2.4. For a given point $\hat{m}\in\hat{\Gamma}$, choose $h\in G$ so that $\xi^{+}(\tau^{+}(\hat{m}))=hP_{1}^{+}$ and $\xi^{-}(\tau^{-}(\hat{m}))=hP_{1}^{-}$ and denote by $L_{h}:G\rightarrow G$ the left translation by $h\in G$. Then consider the tangent spaces
$\mathsf{T}_{hP_{1}^{+}}\mathbb{P}(\mathbb{R}^{d})$ $=\Big\{dL_{h}d\pi^{+}\left(X\right):X\in\bigoplus_{i=2}^{d}\mathbb{R}E_{i1}\Big\}$
$\mathsf{T}_{hP_{1}^{-}}\mathsf{Gr}_{d-1}(\mathbb{R}^{d})$ $=\Big\{dL_{h}d\pi^{-}\left(X\right):X\in\bigoplus_{i=2}^{d}\mathbb{R}E_{1i}\Big\}$
where $E_{ij}$ is the $d\times d$ matrix whose $(i,j)$ entry is $1$ and all the others zero. For $u\in\{0\}\times\mathbb{R}^{d-1}$ we denote by $X_{u}^{+}\in\mathsf{T}_{hP_{1}^{+}}\mathbb{P}(\mathbb{R}^{d})$ and $X_{u}^{-}\in\mathsf{T}_{hP_{1}^{-}}\mathsf{Gr}_{d-1}(\mathbb{R}^{d})$ the tangent vectors
@@
in the fibers of the bundles $\sigma_{\ast}\mathcal{E}^{\pm}\rightarrow\Gamma\backslash\hat{\Gamma}$ over $x=[\hat{m}]_{\Gamma}$ and $(\pi^{+},\pi^{-}):\mathsf{SL}_{d}(\mathbb{R})\rightarrow\mathbb{P}(\mathbb{R}^{d})\times\mathsf{Gr}_{d-1}(\mathbb{R}^{d})$ are the natural projections.
The following lemma shows that when the geodesic flow on $\sigma_{\ast}\mathcal{E}^{-}$ is weakly contracting then the geodesic flow on $\sigma_{\ast}\mathcal{E}^{+}$ is weakly dilating. Recall that $\langle\cdot,\cdot\rangle$ is the standard Euclidean inner product on $\mathbb{R}^{d}$.
@@
For two sequences of positive real numbers $(a_{n})_{n\in\mathbb{N}}$, $(b_{n})_{n\in\mathbb{N}}$ we write $a_{n}\asymp b_{n}$ if there exists $R>0$ such that $R^{-1}a_{n}\leqslant b_{n}\leqslant Ra_{n}$ for every $n\in\mathbb{N}$. We may assume that $\rho(\Gamma)$ is contained in $\mathsf{SL}^{\pm}_{d}(\mathbb{R})$, otherwise we may replace $\rho$ with $\hat{\rho}(\gamma)=|\textup{det}(\rho(\gamma))|^{-1/d}\rho(\gamma)$, $\gamma\in\Gamma$ since $\xi^{\pm}$ are also $\hat{\rho}-$equivariant.
[…]
@@
Notice that for $g\in P_{1}^{\pm}$ we have $dL_{g}\circ d\pi^{\pm}=d\pi^{\pm}\circ\textup{Ad}(g)$ and an elementary calculation shows that
@@
$=dL_{k_{1n}}\left(d\pi^{+}\left(\begin{pmatrix}0&0\\ \frac{1}{\lambda_{n}}A_{n}u&0\end{pmatrix}\right)\right)$.
Similarly, we check that
@@
By the continuity of the family of norms $\big(||\cdot||_{x}\big)_{x\in\Gamma\backslash\hat{\Gamma}}$ and since $k_{1n},k_{2n}\in K$ lie in a compact group, we deduce that
$\left\|\varphi_{t_{n}}\big(X_{u}^{+}\big)\right\|_{\varphi_{t_{n}}(x)}\asymp\frac{\left\|A_{n}u\right\|}{|\lambda_{n}|}\ \ \mathrm{and}\ \ \left\|\varphi_{t_{n}}\big(X_{u}^{-}\big)\right\|_{\varphi_{t_{n}}(x)}\asymp|s_{n}|\cdot\left\|B_{n}^{-t}u\right\|,$
@@
[…]
@@
  1. (i)
[…]
@@
  2. (ii)
[…]
@@
[…]
and hence part (i) follows. The argument for part (ii) is similar. ∎
@@ 4. The Cartan property and the uniform gap summation property
Let $G$ be a real semisimple Lie group of non-compact type, $K$ a maximal compact subgroup of $G$ and $\mu:G\rightarrow\overline{\mathfrak{a}}^{+}$ the associated Cartan projection. The restricted Weyl group of $\mathfrak{a}$ in $\mathfrak{g}$ is the group $W=N_{K}(\mathfrak{a})/Z_{K}(\mathfrak{a})$, where $N_{K}(\mathfrak{a})$ (resp. $Z_{K}(\mathfrak{a})$) is the normalizer (resp. centralizer) of $\mathfrak{a}$ in $K$. The group $W$ is finite, acts simply transitively on the set of Weyl chambers of $\mathfrak{a}$ and contains a unique order two element $w_{0}Z_{K}(\mathfrak{a})\in W$ such that $\textup{Ad}(w_{0})\overline{\mathfrak{a}}^{+}=-\overline{\mathfrak{a}}^{+}$. The element $w_{0}\in K$ defines the opposition involution $^{\star}:\Delta\rightarrow\Delta$ on the set of simple restricted roots $\Delta$ as follows: if $\alpha\in\Delta$ then $\alpha^{\star}\in\Delta$ is the unique root with $\alpha^{\star}(H)=-\alpha(\textup{Ad}(w_{0})H)$ for every $H\in\mathfrak{a}$.
@@
$\lim_{|\gamma|_{\Gamma}\rightarrow\infty}\alpha\big(\mu(\rho(\gamma))\big)=+\infty$
@@
$\Xi_{\theta}^{+}(g):=k_{g}P_{\theta}^{+}\ \ \textup{and}\ \ \Xi_{\theta}^{-}(g):=k_{g}w_{0}P_{\theta}^{-}.$
@@
  1. (1)
    Suppose that $\rho$ admits a continuous $\rho$-equivariant map $\xi^{+}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$. The map $\xi^{+}$ satisfies the Cartan property if for any $x\in\partial_{\infty}\Gamma$ and every infinite sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of elements of $\Gamma$ with $\lim_{n}\gamma_{n}=x$,
@@
  2. (2)
    Suppose that $\rho$ admits a continuous $\rho$-equivariant map $\xi^{-}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{-}$. The map $\xi^{-}$ satisfies the Cartan property if the map $\xi^{\ast}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta^{\star}}^{+}$ satisfies the Cartan property. In other words, for every $x\in\partial_{\infty}\Gamma$ and every infinite sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of elements of $\Gamma$ with $\lim_{n}\gamma_{n}=x$,
    $\xi^{-}(x)=\lim_{n\rightarrow\infty}\Xi^{-}_{\theta}\big(\rho(\gamma_{n})\big).$
Remarks 4.2.
(i) Let $\rho:\Gamma\rightarrow G$ be a $P_{\theta}$-divergent representation. The Cartan property for a continuous $\rho$-equivariant map $\xi^{+}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ (resp. $\xi^{-}$) is independent of the choice of the Cartan decomposition of $G$. This follows by the fact that all Cartan subspaces of $G$ are conjugate under the adjoint action of $G$ and the second part of [24, Cor. 5.9].
(ii) For a $\theta$-divergent sequence $(g_{n})_{n\in\mathbb{N}}\subset G$, written in the Cartan decomposition of $G$ as $g_{n}=k_{n}\exp(\mu(g_{n}))k_{n}^{\prime}$, the condition of $\lim_{n}k_{n}P_{\theta}^{+}=\lim_{n}\Xi_{\theta}^{+}(g_{n})$ to exist, implies that $(g_{n})_{n\in\mathbb{N}}$ $\tau_{\textup{mod}}$-flag converges to $x$, in the definition of Kapovich–Leeb–Porti [29, Subsec. 4.5].
Given a $P_{\theta}$-divergent representation $\rho$ and a $\rho$-equivariant continuous map $\xi:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ with the Cartan property, the map $\Xi_{\theta}^{+}:\Gamma\rightarrow G/P_{\theta}^{+}$, $\gamma\mapsto\Xi_{\theta}^{+}(\rho(\gamma))$ extends continuously to a map $\Gamma\cup\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ restricting to $\xi$ on $\partial_{\infty}\Gamma$.
@@
Suppose that $\rho,\Gamma,G$ and $\theta$ are defined as in Definition 4.1 and let $\xi:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ be a continuous $\rho$-equivariant map. Suppose that $\tau_{\theta}:G\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ is an irreducible $\theta$-compatible proximal representation as in Proposition 2.4 so that $\tau_{\theta}(P_{\theta}^{+})$ stabilizes a line in $\mathbb{R}^{d}$ and induces a $\tau_{\theta}$-equivariant embedding $\iota^{+}:G/P_{\theta}^{+}\xhookrightarrow{}\mathbb{P}(\mathbb{R}^{d})$. The map $\xi$ satisfies the Cartan property if and only if $\iota^{+}\circ\xi$ satisfies the Cartan property.
We need the following estimates which help us verify, in several cases, the Cartan property of limit maps into the homogeneous spaces $G/P_{\theta}^{+}$ and $G/P_{\theta}^{-}$. The second part of the following proposition has been established in [9, Lem. A4] and [24, Lem. 5.8 (i)] but for completeness we give a short proof.
@@
Let $G$ be a real semisimple Lie group, $\theta\subset\Delta$ a subset of simple restricted roots of $G$ and $\tau_{\theta}:G\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ an irreducible, $\theta$-proximal representation such that $\tau_{\theta}(P_{\theta}^{+})$ stabilizes the line $e_{1}]$ in $\mathbb{P}(\mathbb{R}^{d})$. Let $\chi_{\tau_{\theta}}\in\mathfrak{a}^{\ast}$ be the highest weight of $\tau_{\theta}$ and $g,r\in G$.
  1. (i)
    If $g$ is $P_{\theta}$-proximal in $G/P_{\theta}^{+}$ with attracting fixed point $x_{g}^{+}\in G/P_{\theta}^{+}$, then
    $d_{G/P_{\theta}^{+}}\big(x_{g}^{+},\Xi_{\theta}^{+}(g)\big)$ $\leqslant\exp\Big(-\min_{\alpha\in\theta}\alpha(\mu(g))+\chi_{\tau_{\theta}}(\mu(g)-\lambda(g))\Big)$
  2. (ii)
    If $\min_{\alpha\in\theta}\langle\alpha,\mu(g)\rangle>0$ and $\min_{\alpha\in\theta}\langle\alpha,\mu(gr)\rangle>0$, then
    $d_{G/P_{\theta}^{+}}\big(\Xi_{\theta}^{+}(gr),\Xi_{\theta}^{+}(g)\big)\leqslant C_{d,r}\exp\Big(-\min_{\alpha\in\theta}\alpha(\mu(g))\Big)$
    where $C_{d,r}=\sigma_{1}(\tau_{\theta}(r))\sigma_{1}(\tau_{\theta}(r^{-1}))\sqrt{d-1}$}.
@@
$g=h\begin{pmatrix}\ell^{\prime}_{1}(g)&0\\ 0&A_{g}\\ \end{pmatrix}h^{-1}=k_{g}\exp(\mu(g))k_{g}^{\prime},\ \left|\ell_{1}^{\prime}(g)\right|=\ell_{1}(g).$
@@
$h\begin{pmatrix}\ell_{1}^{\prime}(g)&0\\ 0&A_{g}\end{pmatrix}h^{-1}=w_{1}\begin{pmatrix}\ell_{1}^{\prime}(g)&\ast\\ 0&\ast\end{pmatrix}w_{1}^{-1}$
hence $k_{g}^{-1}w_{1}\begin{pmatrix}[0.7]\ell_{1}^{\prime}(g)&\ast\\ 0&\ast\end{pmatrix}$=} $\exp(\mu(g))k_{g}^{\prime}w_{1}$} and $\left|\left\langle k_{g}^{-1}w_{1}e_{1},e_{i}\right\rangle\right|=\frac{\sigma_{i}(g)}{\ell_{1}(g)}\left|\left\langle k_{g}^{\prime}w_{1}e_{1},e_{i}\right\rangle\right|$ for $i>1$. Therefore,
@@
Since $\min_{\alpha\in\theta}\alpha\big(\mu(\rho(\gamma))\big)=\log\frac{\sigma_{1}(\tau_{\theta}(g))}{\sigma_{2}(\tau_{\theta}(g))}$ and $\chi_{\tau_{\theta}}(\lambda(g))=\log\ell_{1}(\tau_{\theta}(g))$ for $g\in G$, part (i) follows.